# CITER: Collaborative Inference for Efficient Large Language Model Decoding with Token-Level Routing

### v4

@@ 3.2 Overall Performance
We conduct extensive experiments to assess the performance of CITER across all benchmark datasets, comparing it against baseline methods. The results are presented in Figure 2. Clearly, all token-level routing methods, including CITER and Co-LLM, significantly outperform the query-level routing method, RouteLLM, across all datasets, particularly on the Commonsense QA and GSM8k datasets, reducing up to 30% inference costs while maintaining the same accuracy or achieving up to 12% higher accuracy with the same cost. This emphasizes the effectiveness of token-level routing, which provides enhanced flexibility in reducing inference costs while preserving response quality. Notably, Speculative Decoding does reduce inference costs on some multiple-choice datasets. However, its verification mechanism requires the small model to produce outputs identical to those of the large model to maintain lossless output quality, which is overly stringent and limits the potential for further cost reduction on complex cases. As a result, on mathematical datasets, the acceptance rate of candidate tokens proposed by the small model is extremely low, leading to higher inference costs than simply using the large model alone, which is unacceptable. In addition, its , Furthermore, CITER consistently surpasses Co-LLM, achieving comparable accuracy with up to 27% fewer inference costs or delivering up to a 17% improvement in accuracy with the same cost. These findings demonstrate the success of our framework in accelerating LLM inference. This outcome is expected, as Co-LLM does not consider long-term information during the router training phase, which is crucial for token-level routing. In the following section, we present experiments to further demonstrate the importance of incorporating long-term information in router training.
@@
Figure 5: Accuracy vs. inference costs curve of CITER with router over the first three iterations on the commonsense QA datasets. Points closer to the top-left corner indicate better acceleration performance.
@@ 3.4 Analysis of Iterative Training Process
To highlight the importance of the iterative training process, we present the performance curve of CITER with the router over the first three iterations on the Commonsense QA dataset. As shown in Figure 5, the results demonstrate a clear improvement in performance in the first two iterations. In the second iteration, CITER reduces $\sim 5\%$ inference costs while maintaining the same accuracy or achieves $2\sim 3\%$ higher accuracy with the same cost compared to the first. This improvement underscores the effectiveness of our proposed iterative training process. Moreover, the performance curve of the third iteration closely follows that of the second, indicating that the router has already converged by the second iteration. The rapid convergence of the router emphasizes the robustness of our training strategy, suggesting that optimal performance can be achieved without excessive inference costs.
3.5 Compatibility Analysis
@@
Additionally, we conduct experiments with Llama3.1 series models to demonstrate the compatibility of our framework. Specifically, in this section, we leverage the Llama3.1-70B model as the LLM and the Llama3.1-8B model as the SLM. The results are illustrated in Figure 6. Similarly to the results with Qwen series, CITER consistently outperforms all other baseline methods, achieving comparable accuracy with up to 32% fewer inference costs or providing up to a 5% improvement in accuracy with the same cost, compared to Co-LLM, the best baseline method. This result further demonstrates the effectiveness of our framework and additionally highlights the compatibility of our framework with different series of models.
@@ 3.6 Analysis of the Impact of SLM Model Size
3.7 Case Study Analysis on the Router
@@ 4 Related Work
Previous routing methods (Jang et al., 2023; Chronopoulou et al., 2023; Diao et al., 2023; Lu et al., 2023; Cheng et al., 2024; Lu et al., 2024; Chen et al., 2023b; Wang et al., 2024b; Srivatsa et al., 2024; Stripelis et al., 2024) for efficient inference mainly focus on routing entire user queries to different models for generation. For example, Routoo (Mohammadshahi et al., 2024) proposes a performance predictor and a cost-aware decoder to route between LLMs, considering both performance and resource constraints; Hybird LLM (Ding et al., 2024) proposes a probabilistic router to select LLM backend for each query; RouteLLM (Ong et al., 2024) formulates the routing problem as a classification problem and employs a data augmentation framework to significantly expand the dataset used for training the router. However, as highlighted in Section 1, routing at the query-level granularity may lead to suboptimal performance, as non-critical tokens in complex queries may be generated inefficiently, while critical tokens in simple queries may suffer from inaccuracy. In contrast, token-level routing methods offer more fine-grained control over the routing process, improving both inference costs and the quality of the generated response.
@@
Unlike query-level routing methods, previous token-level routing methods (Pfeiffer et al., 2021; Belofsky, 2023; Muqeeth et al., 2024; Wang et al., 2024a; Wu et al., 2024; Xu et al., 2024) mainly focus on routing input tokens to different specialized experts to enhance performance without considering the inference costs. For example, Arrow (Ostapenko et al., 2024) build a mixture-of-experts (MoE) architecture with multiple LoRAs, dynamically routing inputs to different LoRAs during inference. Among these methods, Co-LLM (Shen et al., 2024) is the most relevant to our framework CITER, introducing a router to route tokens to models of different sizes. However, Co-LLM only considers the current outputs from SLM and LLM when generating ground truth labels to train the router. This may lead to suboptimal performance since the influence of current decisions on future tokens is not considered. Moreover, similar to other token-level routing methods, Co-LLM focuses on enhanced response quality without taking the inference costs of the inference process into account. In contrast, our CITER framework considers both the current token and the future impact of each decision, enabling more accurate and efficient routing.
@@ Appendix A Acknowledgment
This work was supported in part by the National Institutes of Health (NIH) under grants R01AG085581 and R01AG079291.

### v6

@@ 3.2 Overall Performance
We conduct extensive experiments to assess the performance of CITER across all benchmark datasets, comparing it against baseline methods. The results are presented in Figure 2. Clearly, all token-level routing methods, including CITER and Co-LLM, significantly outperform the query-level routing method, RouteLLM, across all datasets, particularly on the Commonsense QA and GSM8k datasets, reducing up to 30% inference costs while maintaining the same accuracy or achieving up to 12% higher accuracy with the same cost. This emphasizes the effectiveness of token-level routing, which provides enhanced flexibility in reducing inference costs while preserving response quality. Notably, Speculative Decoding does reduce inference costs on some multiple-choice datasets. However, its verification mechanism requires the small model to produce outputs identical to those of the large model to maintain lossless output quality, which is overly stringent and limits the potential for further cost reduction on complex cases. As a result, on mathematical datasets, the acceptance rate of candidate tokens proposed by the small model is extremely low, leading to higher inference costs than simply using the large model alone, which is unacceptable. Furthermore, CITER consistently surpasses Co-LLM, achieving comparable accuracy with up to 27% fewer inference costs or delivering up to a 17% improvement in accuracy with the same cost. These findings demonstrate the success of our framework in accelerating LLM inference. This outcome is expected, as Co-LLM does not consider long-term information during the router training phase, which is crucial for token-level routing. In the following section, we present experiments to further demonstrate the importance of incorporating long-term information in router training.
@@
Figure 5: Accuracy vs. inference costs of CITER with router over the first three iterations on the commonsense QA datasets. Points closer to the top-left corner indicate better acceleration performance.
@@ 3.4 Analysis of Iterative Training Process
To highlight the importance of the iterative training process, we present the performance curve of CITER with the router over the first three iterations on the Commonsense QA dataset. As shown in Figure 5, the results demonstrate a clear improvement in performance in the first two iterations. In the second iteration, CITER reduces $\sim 5\%$ inference costs while maintaining the same accuracy or achieves $2\sim 3\%$ higher accuracy with the same cost compared to the first. This improvement underscores the effectiveness of our proposed iterative training process. Moreover, the performance curve of the third iteration closely follows that of the second, indicating that the router has already converged by the second iteration. The rapid convergence of the router emphasizes the robustness of our training strategy, suggesting that optimal performance can be achieved without excessive costs.
3.5 Results on Different Model Families
@@
Additionally, we conduct experiments with Llama3.1 series models to demonstrate the compatibility of our framework. Specifically, we leverage the Llama3.1-70B model as the LLM and the Llama3.1-8B model as the SLM. The results are illustrated in Figure 6. Similarly to the results with Qwen series, CITER consistently outperforms all other baseline methods, achieving comparable accuracy with up to 32% fewer inference costs or providing up to a 5% improvement in accuracy with the same cost, compared to Co-LLM, the best baseline method. This result further demonstrates the effectiveness of our framework and additionally highlights the compatibility of our framework with different series of models.
Figure 7: Accuracy vs. latency curve of CITER and Speculative Decoding.
@@ 3.6 Analysis of the Impact of SLM Model Size
3.7 Latency Analysis
In addition, we also evaluate the wall-clock latency of CITER against Speculative Decoding. The results are presented in Figure 7. The plot shows that CITER offers a flexible trade-off between latency and accuracy. For instance, CITER achieves an accuracy of 80.8% with a latency of only 4 seconds. While Speculative Decoding reaches a slightly higher accuracy of 87.0%, it requires a significantly longer latency of 10.9 seconds. CITER can also achieve 85.6% accuracy which is very similar to Speculative Decoding with the same latency. In addition, key advantage of CITER lies in the flexibility to operate at much lower latency points with only a minor compromise in accuracy, a feature not available with Speculative Decoding. This demonstrates the superior efficiency and adaptability of our framework.
3.8 Qualitative Analysis on the Router
@@ 4 Related Work
Previous routing methods (Jang et al., 2023; Chronopoulou et al., 2023; Diao et al., 2023; Lu et al., 2023; Cheng et al., 2024; Lu et al., 2024; Chen et al., 2023b; Wang et al., 2024b; Srivatsa et al., 2024; Stripelis et al., 2024) for efficient inference mainly focus on routing entire user queries to different models for generation. For example, Routoo (Mohammadshahi et al., 2024) proposes a performance predictor and a cost-aware decoder to route between LLMs, considering both performance and resource constraints; Hybird LLM (Ding et al., 2024) proposes a probabilistic router to select LLM backend for each query; RouteLLM (Ong et al., 2024) formulates the routing problem as a classification problem and employs a data augmentation framework to significantly expand the dataset used for training the router; Gupta et al. (2024) first leverages the small model (SLM) to generate the entire sequence, and then the decision to defer to the large model (LLM) is made based on the uncertainty (e.g., entropy or confidence derived from logits) during generation;FrugalGPT Chen et al. (2023b) formulates the routing problem as a constrained optimization problem, where the final generated quality is maximized under a budget or inference cost constraint. However, as highlighted in Section 1, routing at the query-level granularity may lead to suboptimal performance, as non-critical tokens in complex queries may be generated inefficiently, while critical tokens in simple queries may suffer from inaccuracy. In contrast, token-level routing methods offer more fine-grained control over the routing process, improving both inference costs and the quality of the generated response.
@@
Unlike query-level routing methods, previous token-level routing methods (Pfeiffer et al., 2021; Belofsky, 2023; Muqeeth et al., 2024; Wang et al., 2024a; Wu et al., 2024; Xu et al., 2024) mainly focus on routing input tokens to different specialized experts to enhance performance without considering the inference costs. For example, Arrow (Ostapenko et al., 2024) builds a mixture-of-experts (MoE) architecture with multiple LoRAs, dynamically routing inputs to different LoRAs during inference. Similarly, Branch-Train-MiX (Sukhbaatar et al., 2024) fine-tunes LLMs on different domains from a seed LLM, creating specialized experts to form an MoE framework. Besides these methods, Narasimhan et al. (2024) introduces token-level cascading and tackles the challenge of implementing deferral rules by leveraging speculative decoding, relying exclusively on the logits output by the model to determine when to defer to the LLM. Similarly, Co-LLM (Shen et al., 2024) introduces a router to route tokens to models of different sizes. However, they only consider the current outputs from SLM and LLM when generating ground truth labels to make the router decisions. This may lead to suboptimal performance since the influence of current decisions on future tokens is not considered. Moreover, similar to other token-level routing methods, they focus on enhanced response quality without taking the inference costs of the inference process into account. In contrast, our CITER framework considers both the current token and the future impact of each decision, enabling more accurate and efficient routing.
@@ Appendix A Acknowledgment
This work was supported in part by the National Institutes of Health (NIH) under grants R01AG079291, R01AG085581 and R01HL165061, National Science Foundation (NSF) under grants CNS2414087, NSF BCS2040381, NSF IIS2123952, NSF IIS1955532, NSF IIS2123952, NSF IIS2311990, and the Semiconductor Research Corporation (SRC) AIHW award 2024AH3210.